OpenStudy (anonymous):

working alone takes mark 12 hrs.Beth does it in 10 hours.find hw long it would take them if they worked together

OpenStudy (amistre64):

M = 1/12; B = 1/10 M + B = 22/120. 120 /22 = 5.45 hours round about??

OpenStudy (amistre64):

11/60 55/60 = 5 left 11/5 = 2.2 pqrts of an hour ....5 hours and 27 minutes? best guess

OpenStudy (anonymous):

x x --- + ---- = 1 12 10 multiply each term by 60 (common denominator) 60x 60x --- + ---- = 1(60) 12 10 reduce 5x + 6x = 60 11x = 60 x = 5 5/11 hours